OpenStudy (anonymous):

hey @DanJS and @Nnesha Solve. –24 = –31 + w A. –55 B. –7 C. 7 D. 55

OpenStudy (danjs):

just add 31 to both sides of the = -24 + 31 = w

Nnesha (nnesha):

add 31 both side

OpenStudy (phebe):

-24-(-31+w)=0 Step by step solution : Step 1 : Equation at the end of step 1 : Step 2 : Solve 7-w = 0 Solving a Single Variable Equation : 2.1 Solve : -w+7 = 0 Subtract 7 from both sides of the equation : -w = -7 Multiply both sides of the equation by (-1) : w = 7 ANSWER IS: w = 7

Nnesha (nnesha):

when you move number or variables from left to right or right to left you must have to change sign so if 31 is negative opposite of negative is positive so add 31 both side

OpenStudy (phebe):

so C is the answer

OpenStudy (anonymous):

c

Nnesha (nnesha):

yes